### v1

# Decentralized Autoregressive Generation

Abstract
We present a theoretical analysis of decentralization of autoregressive generation. We define the Decentralized Discrete Flow Matching objective, by expressing probability generating velocity as a linear combination of expert flows. We also conduct experiments demonstrating the equivalence between decentralized and centralized training settings for multimodal language models across diverse set of benchmarks. Specifically, we compare two distinct paradigms: LLaVA and InternVL 2.5-1B, which uses a fixed CLIP vision encoder and performs full-parameter fine-tuning (ViT+MLP+LLM) during the instruction tuning stage.
1 Introduction
The increase in performance of Multimodal Large Language Models (MLLMs) requires a massive increase in parameter counts (kaplan2020scalinglawsneurallanguage), which, in turn, requires enhanced infrastructure: training requires centralized clusters of thousands of GPUs with extensive gradient synchronization.
These developments present two critical challenges. First, training LLMs is inaccessible to academic institutions and small startups due to limited infrastructure. Second, the system is susceptible to individual node failures, which can disrupt the entire process (grattafiori2024llama3herdmodels). Existing mitigation strategies include aggressive checkpointing and related techniques.
A potential solution is a decentralized training approach, where multiple expert models are trained independently on disjoint subsets of the dataset, with only a few activated during inference time. This significantly reduces communication between nodes and enables academic institutions to collaborate. Recent studies (li2022branchtrainmergeembarrassinglyparalleltraining; gururangan2023scalingexpertlanguagemodels; no-need-to-talk) empirically demonstrate the efficiency of this approach for unimodal language modeling training in specific settings; however, existing work lacks a strict theoretical guarantee that decentralized training is equivalent to the centralized paradigm, making decentralized training a black box. Moreover, decentralized training of multimodal large language models has not yet been empirically explored.
mcallister2025decentralizeddiffusionmodels defined the Decentralized Flow Matching objective for Continuous Flow Matching (lipman2023flow) and demonstrated the equivalence between centralized and decentralized training of flow-matching models for image generation. At the same time, gat2024discrete introduced the Discrete Flow Matching theoretical framework, extending Continuous Flow Matching (lipman2023flow) to a discrete data domain. Inspired by these advancements, we derive the theoretical analysis of decentralized autoregressive generation.
Our contributions are the following:
  • We extend the Discrete Flow Matching paradigm (gat2024discrete) to the discrete time domain, making the necessary modifications and defining the corresponding conditions.
  • Similar to mcallister2025decentralizeddiffusionmodels, we define the Decentralized Discrete Flow Matching objective and show that the probability-generating velocity is a linear combination of expert flows.
  • We show that autoregressive generation is a special degenerate case of the Discrete Flow Matching theoretical paradigm. We present the corresponding probability path and probability generating velocity, implying that the Decentralized Discrete Flow Matching objective can be applied to autoregressive generation.
  • We conduct experiments demonstrating the equivalence between decentralized and centralized training for multimodal language models in two distinct settings: LLaVA (liu2024improvedbaselinesvisualinstruction) and InternVL (chen2025expandingperformanceboundariesopensource).
2 Background
In this section, we briefly reintroduce Discrete Flow Matching theoretical framework.
2.1 Continuous Time Discrete Flow Matching
Discrete Flow Matching (gat2024discrete) is a theoretical framework that describes discrete generative modeling processes in terms similar to Continuous Flow Matching (lipman2023flow). Namely, having vocabulary or tokens $[d]$, source distribution $X_{0}$ with PMF $p(x)$ of token sequences $x=(x^{1},\ldots,x^{N})\in[d]^{N}$ and target distribution $X_{1}$ with PMF $q(x)$, we define the probability path $p_{t}(x),t\in[0,1]$ s.t. $p_{0}(x)=p(x)$, $p_{1}(x)=q(x)$ as follows:
$p_{t}(x)=\sum_{(x_{0},x_{1})\in[d]^{N\times N}}p_{t}(x|x_{0},x_{1})\pi(x_{0},x_{1})$ (1)
$p_{t}(x|x_{0},x_{1})=\prod_{i=1,\ldots N}p_{t}(x^{i}|x_{0},x_{1})$ (2)
where $\pi(x_{0},x_{1})$ is a joint PMF of $(X_{0},X_{1})$, and $p_{t}(x^{i}|x_{0},x_{1})$ is a conditional marginal probability path for the token position $i$.
The $p_{t}(x|x_{0},x_{1})$ should satisfy the following conditions:
$p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)=\begin{cases}1,&x=x_{0}\\ 0,&x\neq x_{0}\end{cases}$ (3)
$p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)=\begin{cases}1,&x=x_{1}\\ 0,&x\neq x_{1}\end{cases}$ (4)
In the framework, in order for probability paths to be tractable, we consider only the probability paths that are represented by a convex sum of conditional probabilities $w^{j}(x^{i}|x_{0},x_{1})$ with coefficients provided by some scheduler $\kappa^{i,j}_{t}$:
$p_{t}(x^{i}|x_{0},x_{1})=\sum_{j}\kappa^{i,j}_{t}w^{j}(x^{i}|x_{0},x_{1})$ (5)
$\sum_{j}\kappa^{i,j}_{t}=1,\quad\kappa_{t}^{i,j}\geq 0$ (6)
We also define probability generating velocity $u^{i}_{t}(x,z),i\in[N]$. We sample our token sequences by the following rule (for each token position independently):
$X_{t+h}^{i}\sim\delta_{X_{t}}(\cdot)+hu_{t}^{i}(\cdot,X_{t}),\quad h>0$ (7)
We say that $u_{t}$ generates the probability path $p_{t}(x)$ if $X_{t+h}$, sampled by the rule 7, satisfies:
$X_{t+h}\sim p_{t+h}+o(h)$ (8)
Probability generating velocity $u_{t}$ can be constructed through conditional probability generating velocities $u^{i}_{t}(x|x_{0},x_{1})$. In fact, the following theorem holds:
Theorem 1
Probability generating velocity $u_{t}^{i}(x^{i})$ defined as follows:
$u_{t}^{i}(x^{i},z)=\sum_{(x_{0},x_{1})\in[d]^{N\times N}}u_{t}^{i}(x^{i},z|x_{0},x_{1})\frac{p_{t}(z|x_{0},x_{1})\pi(x_{0},x_{1})}{p_{t}(z)}$ (9)
generates the probability path $p_{t}(x)$
This is proven by using the Continuity Equation:
$\dot{p_{t}}(x)+div_{x}(p_{t}u_{t})=0$ (10)
The discrete divergence of $p_{t}u_{t}$ is defined in the following way:
$div_{x}(v)=\sum_{(x,z)\in[d]^{N\times N}}v(z,x)-v(x,z)$ (11)
where $v(z,x)=p_{t}(x)u_{t}^{i}(z^{i},x)$, $v(x,z)=p_{t}(z)u_{t}^{i}(x^{i},z)$ if $z$ differs from $x$ only in the token position $i$, $v(x,x)=p_{t}(x)\sum_{i=1}^{N}u_{t}^{i}(x^{i},x)$ and $v(z,x)=v(x,z)=0$ otherwise.
The divergence of $p_{t}u_{t}$ can be shown to be equal to:
$div_{x}(p_{t}u_{t})=-\sum_{z\in[d]^{N}}p_{t}(z)\sum_{i=1}^{N}\delta_{z}(x^{\bar{i}})u_{t}(x^{i},z)$ (12)
where $\delta_{z}(x^{\bar{i}})=\prod_{j\neq i}\delta_{z}(x^{j})$.
As shown in (gat2024discrete), if the continuity equation holds, then the probability generating velocity generates a probability path as defined in (7-8).
3 Theoretical Analysis
In this section, we formulate autoregressive generation as a special case of Discrete Flow Matching. This allows us to define the Discrete Decentralized Flow Matching objective and apply it to AR generation, theoretically showing the equivalence of decentralized and centralized training paradigms. Since AR generation is done in discrete time steps, we first extend DFM to the discrete time domain, introducing necessary modifications. Next, we define the probability path and the probability generating velocity for AR sampling. Since checking that probability generating velocity indeed generates probability path is reduced to checking the Continuity Equation (CE), we define a necessary and sufficient 1-sparsity condition that discrete probability generating velocity must satisfy so that CE implies generation. Finally, we define the Discrete Decentralized Flow Matching objective, which is automatically applicable to AR, thereby proving the equivalence of the decentralized and centralized settings.
3.1 Discrete Time Discrete Flow Matching
In this subsection, we define Discrete Flow Matching in the discrete time domain.
We begin by transitioning to a discrete-time domain, $t\in\{0,1,\ldots,n\}$, and correspondingly redefining the terms. In this setting, the sampling rule 7 becomes:
$X_{t+1}\sim\delta_{X_{t}}(\cdot)+u_{t}(\cdot,X_{t})\quad t\in\{0,1,\ldots,n-1\}$ (13)
where probability generating velocity $u_{t}$ is now defined in discrete timesteps. $u_{t}$ now generates $p_{t}(x)$ if $X_{t+1}$ sampled by the rule 12 satisfies the following:
$X_{t+1}\sim p_{t+1}$ (14)
Here, in order to define proper PMF, it is necessary and sufficient that probability velocity $u_{t}$ satisfies the following conditions:
$\sum_{x^{i}\in[d]}u_{t}^{i}(x^{i}|z)=0$ (15)
$-1\leq u_{t}^{i}(z^{i}|z)\leq 0,\quad 0\leq u_{t}^{i}(x^{i}|z)\leq 1\quad\text{for}\;x^{i}\neq z^{i}$ (16)
The continuity equation now takes the form:
$p_{t+1}(x)-p_{t}(x)+div_{x}(p_{t}u_{t})=0$ (17)
3.2 Autoregressive sampling
To define autoregressive sampling, we define the source and target distribution as a special case of C-coupling from (gat2024discrete):
$(X_{0},X_{1})\sim({\mathbb{I}}\odot X_{1}+(1-{\mathbb{I}})\odot(\mathbf{m},\mathbf{m},\ldots,\mathbf{m}),X_{1})$ (18)
where $X_{1}\sim q$, $\mathbf{m}$ is a mask token, ${\mathbb{I}}\in\{0,1\}^{N}$ is an indicator vector of the form $(1,1,\dots,1,0,\dots,0)$ which has all the ones before some token position and all zeros after.
Next, the condition probability path for a pair $X_{0},X_{1}$ s.t. $X_{0}$ has $P$ tokens revealed at timestep 0 will take the form:
$p_{t}(x^{i}|x_{0},x_{1})=\kappa_{t}^{i}\delta_{X_{1}}(x^{i})+(1-\kappa_{t}^{i})\delta_{X_{0}}(x^{i})$ (19)
$\kappa_{t}^{i}=\begin{cases}0,&t<P+i\\ 1,&t\geq P+i\end{cases}$ (20)
This definition of conditional probability path and scheduler simply means that after timestep $t$ exactly $t+P$ tokens are revealed.
Note that at any time step $t$, $p_{t}(x|x_{0},x_{1})$ is a degenerate distribution with a single result $x_{t}$:
$p_{t}(x|x_{0},x_{1})=\delta_{x_{t}}(x)$
Now, define the conditional generating velocity $u_{t}^{i}(x^{i},z|x_{0},x_{1})$ as:
$u_{t}^{i}(x^{i},x_{t}|x_{0},x_{1})=\begin{cases}\delta_{x_{t+1}}(x^{i})-\delta_{x_{t}}(x^{i}),&z=x_{t}\\ 0,&z\neq x_{t}\end{cases}$ (21)
From the above formula it follows that for a fixed time step $t$ $u_{t}(x^{i},x_{t}|x_{0},x_{1})=0$ for any token position except possibly $i=P+t+1$. This property is necessary and sufficient to show that the continuity equation implies the generation of probability path by probability generation velocity in the discrete time domain.
Next, in order to show that defined conditional probability generating velocity generates conditional probability path, we need to 1) Show that continuity equation holds 2) Show that if the continuity equation holds, then generation follows.
First, check that the continuity equation holds:
$p_{t+1}(x|x_{0},x_{1})-p_{t}(x|x_{0},x_{1})+div_{x}(p_{t}u_{t})=0$ (22)
Since $p_{t}(x|x_{0},x_{1})$ has only one outcome $x_{t}$ and $u_{t}^{i}=0$ for any $i\neq P+t+1$, the divergence formula 12 simplifies to the following:
$div(p_{t}u_{t})=-\delta_{x_{t}}(x^{\overline{P+t+1}})u_{t}(x^{P+t+1},x_{t}|x_{0},x_{1})$ (23)
Substituting 21, we get:
$p_{t+1}(x|x_{0},x_{1})-p_{t}(x|x_{0},x_{1})+div_{x}(p_{t}u_{t})$
$=\delta_{x_{t+1}}(x)-\delta_{x_{t}}(x)-\delta_{x_{t}}(x^{\overline{P+t+1}})u_{t}(x^{P+t+1},x_{t}|x_{0},x_{1})$
$=\delta_{x_{t+1}}(x)-\delta_{x_{t}}(x)-\delta_{x_{t}}(x^{\overline{P+t+1}})(\delta_{x_{t+1}}(x^{P+t+1})-\delta_{x_{t}}(x^{P+t+1}))$
$=\delta_{x_{t+1}}(x)-\delta_{x_{t}}(x)-\delta_{x_{t+1}}(x)+\delta_{x_{t}}(x)$
$=0$
Now, once we verified that the continuity equation holds, we show that satisfying the continuity equation implies generation. For that, we require a 1-sparsity property of generating probability velocity: for a fixed timestep $t$ $u_{t}^{i}=0$ for any token position $i$ except possibly one, denoted by $j=j(t)$ depending only on a timestep $t$. As discussed above, velocity defined in (21) satisfies it. Then the PMF of $X_{t+1}$ sampled by 14 is:
$p_{X_{t+1}}(x)=$ $\sum_{z\in[d]^{N}}p_{t}(z)\prod_{i=1}^{N}\bigl(\delta_{z}(x^{i})+u_{t}^{i}(x^{i},z)\bigr)$
$=\sum_{z\in[d]^{N}}p_{t}(z)\Bigl(\delta_{z}(x)+\delta_{z}(x^{\bar{j}})\,u_{t}^{j}(x^{j},z)\Bigr)$
$=p_{t}(x)+\sum_{z\in[d]^{N}}p_{t}(z)\,\delta_{z}(x^{\bar{j}})\,u_{t}^{j}(x^{j},z)$
$=p_{t}(x)+\operatorname{div}_{x}\bigl(p_{t}u_{t}\bigr)=p_{t}(x)+\left(p_{t+1}(x)-p_{t}(x)\right)=p_{t+1}(x)$
We defined autoregressive sampling as an instance of discrete flow matching in a discrete time domain. We proved that for probability generating velocities that are non-zero at only one token position for any fixed timestep satisfying continuity equation implies generation. We showed that the autoregressive probability path defined in 19 and the velocity defined in 21 satisfy the continuity equation and therefore define a valid probability path and generating velocity.
3.3 Decentralization
The above formalization automatically makes Theorem 1 true for autoregressive sampling. Now, similar to (mcallister2025decentralizeddiffusionmodels), if we define a target distribution $X_{1}$ as a union of disjoint clusters $S_{k},\quad k\in\{1,2,\dots,K\}$, we can rewrite Equation 24 as follows:
$u_{t}^{i}(x^{i},z)=\sum_{(x_{0},x_{1})\in[d]^{N\times N}}u_{t}^{i}(x^{i},z|x_{0},x_{1})\frac{p_{t}(z|x_{0},x_{1})\pi(x_{0},x_{1})}{p_{t}(z)}$ (24)
$=\sum_{k=1}^{K}\frac{1}{p_{t}(z)}\sum_{(x_{0},x_{1})\in S_{k}}u_{t}^{i}(x^{i},z|x_{0},x_{1})p_{t}(z|x_{0},x_{1})\pi(x_{0},x_{1})$ (25)
$=\sum_{k=1}^{K}\frac{p_{t}(z|S_{k})}{p_{t}(z)}\sum_{(x_{0},x_{1})\in S_{k}}u_{t}^{i}(x^{i},z|x_{0},x_{1})\frac{p_{t}(z|x_{0},x_{1})\pi(x_{0},x_{1})}{p_{t}(z|S_{k})}$ (26)
$=\sum_{k=1}^{K}\frac{p_{t}(S_{k}|z)}{p_{t}(S_{k})}u_{t}^{i}(x,z|S_{k})$ (27)
The inner sum is called an expert flow, and the outer sum is called a router. The equation above simply means that theoretically, the overall generating probability velocity is represented as a weighted sum of expert flows.
If we assume that data is sampled from the clusters with equal probabilities and that clusters are convex (i.e. if $X_{1}\in S_{k}$, then $X_{t}\in S_{k}\quad\forall t$), then the above formula simplifies to:
$u_{t}^{i}(x^{i},z)=\frac{1}{K}\sum_{k=1}^{K}p_{t}(S_{k}|z)u_{t}^{i}(x,z|S_{k})$ (28)
4 Method
Since the effectiveness of decentralized discrete flow matching in case of pure LM has already been shown in previous work (no-need-to-talk; gururangan2023scalingexpertlanguagemodels), we focus on Vision Language Models as another instance of autoregressive sampling.
4.1 Expert training
During training time, the image-text pairs dataset is partitioned into $K$ disjoint clusters, and $K$ experts are trained completely independently on their assigned clusters.
4.2 Data partitioning
To obtain training clusters for experts, we use the pretrained vision encoder model (CLIP) to extract image features from unique images in the dataset. We train spherical balanced $k$-means to cluster features into $K$ clusters of equal sizes. Importantly, while this strategy balances the number of unique images per cluster, it does not perfectly balance the number of image-text pairs or total token counts. This discrepancy arises because a single image may correspond to multiple text pairs and caption lengths vary significantly across domains. Consequently, the computational load and required training steps may differ slightly between experts.
4.3 Inference strategy
During inference time, for a sample with image $I$ we set cluster probabilities as:
$p_{t}(S_{k}|z,x)=\frac{\exp(\tau\cdot\cos(x,c_{k}))}{\sum_{j=1}^{K}\exp(\tau\cdot\cos(x,c_{j}))}$ (29)
where $x=f_{\text{CLIP}}(I)$ – CLIP features vector, $\cos(\cdot,\cdot)$ is the cosine similarity, $f_{\text{CLIP}}$ is a feature extractor, $c_{j}$ are cluster centroids, $\tau$ is the temperature parameter. This implies that routing is time independent and agnostic of current token sequence state.
To balance theoretical fidelity with computational efficiency, we employ a top-$k$ strategy. The final output probabilities are top-$k$ filtered and renormalized. For the main experiments, we take $k=1$ to ensure that the ensemble inference computation budget (excluding routing cost) is the same as that of the dense model.
5 Experiments
5.1 LLaVA
In the first part of the experiments, we utilize the LLaVA-1.5 (liu2024improvedbaselinesvisualinstruction) fine-tuning dataset, consisting of 625k image-text samples and 41k text-only samples from diverse sources.
During the data partitioning stage, text-only samples are randomly and equally distributed between the clusters.
We train $K=2$ experts. As an initialization checkpoint for experts, we employ the LLaVA-1.5-7B model after the vision language alignment stage. We effectively perform the visual instruction tuning stage of the experts.
Each expert is trained on the same number of GPU devices as the dense baseline. However, we halve the per-device batch size to ensure that the total number of training steps remains consistent with the original model training schedule. All other hyperparameters remain unchanged.
We utilize CLIP-ViT-L/14@336px as the vision encoder for our routing mechanism. We select this specific version because the LLaVA-1.5 architecture already requires it for visual processing. Consequently, with a shared encoder implementation, the feature extraction step for routing incurs zero additional computational overhead during inference.
5.2 InternVL
In the second phase of our experiments, we utilize a subset of the InternVL 2.5 (chen2025expandingperformanceboundariesopensource) Stage 2 fine-tuning data mixture (see Table 1). We curated datasets representing diverse task types to ensure comprehensive coverage.
Consistent with the first experimental phase 5.1, we train $K=2$ experts. We initialize the experts using the InternVL-2.5-1B-Pretrained checkpoint (post-Stage 1.5 ViT+MLP pretraining) and perform full Stage 2 fine-tuning. For dataset partitioning, we apply the same balanced $k$-means algorithm described previously.
To ensure a fair comparison, both the dense baseline and the experts are trained on an identical number of GPU devices with a per-device batch size of 1. We halve the number of gradient accumulation steps for the expert runs. This ensures that the total number of optimization steps remains comparable to the dense model baseline.
Unlike the standard InternVL training pipeline, we disable dynamic resolution to accelerate training. Additionally, the number of training devices is set to 2, and the context length is capped at 8192 tokens. All other hyperparameters remain unchanged.
We utilize CLIP-ViT-B/16 as a vision encoder for dataset partitioning and routing. This results in additional 5.4% computational overhead during inference compared to dense baseline.
We selected InternVL as an alternative to LLaVA to verify the effectiveness of our expert training strategy across different architectures and training setups. Specifically, this allows us to contrast two distinct paradigms: (1) LLaVA-1.5-7B, which uses a fixed CLIP vision encoder and only fine-tunes the MLP connector and LLM; and (2) InternVL 2.5-1B, which uses an Intern-ViT-300M encoder and performs full-parameter fine-tuning (ViT+MLP+LLM) during the instruction tuning stage.
Task type Datasets
Captioning TextCaps (sidorov2020textcapsdatasetimagecaptioning) , ShareGPT4o (cui2025comprehensive)
General QA VQAv2(balanced_vqa_v2), GQA(hudson2019gqanewdatasetrealworld)
Chart ChartQA(masry2022chartqabenchmarkquestionanswering)
OCR InfoVQA(mathew2021infographicvqa) , TextVQA(singh2019vqamodelsread)
Knowledge KVQA(shahMYP19), A-OKVQA(schwenk2022aokvqabenchmarkvisualquestion)
Document DocVQA(mathew2021docvqadatasetvqadocument)
Grounding RefCOCO/+/g (mao2016generationcomprehensionunambiguousobject; yu2016modelingcontextreferringexpressions)
Science AI2DKembhavi2016ADI, ScienceQAlu2022learn
Conversation ALLaVA(chen2024allavaharnessinggpt4vsynthesizeddata)
Table 1: InternVL-type models fine-tuning data mixture
5.3 Evaluation
We evaluate our LLaVA models across total of 8 benchmarks, used in the liu2024improvedbaselinesvisualinstruction. For InternVL, we evaluate across 14 benchmarks, aiming to assess the models in the general question answering, OCR and other tasks, used in the chen2025expandingperformanceboundariesopensource.
5.4 Results
Overall, the expert ensemble achieves near-parity with the compute-matched dense baseline, with systematic trade-offs. In the LLaVA setting, performance is preserved on core QA benchmarks (+1.49+1.49 on VQAv2; 0.03-0.03 on GQA), while OCR-heavy and distribution-shifted benchmarks degrade (1.53-1.53 on TextVQA; 4.47-4.47 on VisWiz). The decrease on MME (33.32-33.32) is consistent with a modest reduction in broad multi-skill coverage under partition-induced specialization.
Method VQAv2 GQA VizWiz SciQA-IMG TextVQA
Dense baseline 78.50 62.00 50.00 66.80 58.20
2 experts 79.99 61.97 45.53 67.03 56.67
Table 2: LLaVA experts results on academic-task-oriented datasets
Method POPE MME MMBench
adv rand pop en zh
Dense baseline 85.9 87.3 86.1 1510.7 64.3 58.3
2 experts 87.1 85.7 87.3 1477.38 65.03 56.87
Table 3: LLaVA experts results on benchmarks for instruction-following LMMs
In the InternVL setting, the expert ensemble remains close to the compute-matched dense baseline on most general benchmarks, with clear, structured trade-offs. On general QA, results are largely preserved (VQAv2: 0.31-0.31, GQA: +1.01+1.01, SciQA-IMG: +0.25+0.25), while broad multi-skill coverage decreases on MME (55.92-55.92). On OCR/document-style tasks, performance is mixed: AI2D and ChartQA improve slightly (+0.84+0.84, +0.28+0.28), DocVQA is unchanged, but TextVQA drops (4.56-4.56), consistent with reduced global/OCR coverage under partitioning. Notably, visual grounding improves substantially: RefCOCO increases by +6+688 points across splits and RefCOCO+ by +5+577 points, indicating that partition-induced specialization can benefit localization-heavy evaluations.
Method AI2D ChartQA TextVQA DocVQA InfoVQA
Dense baseline 61.66 61.96 55.47 48.49 25.02
2 experts 62.50 62.24 50.91 48.49 24.00
Table 4: InternVL experts results on OCR, chart, and document understanding
Method VQAv2 GQA SciQA-IMG POPE (avg) MME
Dense baseline 72.5 56.35 83.64 86.17 1531.7
2 experts 72.19 57.36 83.89 85.55 1475.78
Table 5: InternVL experts results on general QA, multimodal understanding and hallucination performance
RefCOCO RefCOCO+ RefCOCOg
Method val testA testB val testA testB val test
Dense baseline 67.93 73.75 63.55 57.68 65.98 49.21 63.13 62.8
2 experts 75.47 80.20 71.36 62.46 71.29 56.43 71.94 70.45
Table 6: InternVL experts results on visual grounding
5.5 Ablation Study
For InternVL settings, we also conduct an ablation study over 3 factors: number of experts; the choice of vision encoder for image feature extraction for dataset partitioning and routing; and clustering algorithm.
The impact of number of experts
For the main InternVL experiments, we train $K=2$ experts. To test the impact of number of experts, we also train $K=4$ experts using spherical balanced k-means on CLIP ViT-B-16 features. As shown in Table 7, the $K=4$ ensemble maintains performance comparable to the dense baseline, demonstrating the robustness of the decentralized approach. The slight regression compared to the $K=2$ setting is explainable by data fragmentation, where the further partitioning reduces sample density per expert.
Method VQAv2 TextVQA MME RefCOCO
val testA testB
2 experts 72.19 50.91 1475.78 75.47 80.20 71.36
4 experts 70.96 47.59 1330.41 72.05 78.47 66.79
Table 7: The impact of number of experts
The impact of vision encoder
In the main InternVL experiments, we use CLIP ViT-B-16 as a feature extractor for k-means input for dataset partitioning and routing. To investigate the impact of the vision encoder, we also train $K=2$ experts with balanced k-means algorithm using CLIP ViT-L-14@336px. As shown in Table 8, using larger vision encoder leads to the marginal increase in QA (+0.4 VQAv2) and OCR (+0.81 TextVQA) performance.
Method VQAv2 TextVQA MME RefCOCO
val testA testB
ViT-B-16 72.19 50.91 1475.78 75.47 80.20 71.36
ViT-L-14@336px 72.56 51.72 1450.04 74.74 80.75 71.01
Table 8: The impact of vision encoder
The impact of clustering algorithm
In the main InternVL experiments, we use spherical balanced k-means clustering algorithm for dataset partitioning and routing. To test the impact of the clustering algorithm, we also train $K=2$ experts using 2-stage balanced spherical k-means algorithm inspired by mcallister2025decentralizeddiffusionmodels. Namely, first stage is fine clustering into $k=1024$ clusters in unbalanced manner. Second stage is coarse clustering of fine centroids into $K=2$ coarse clusters in a balanced manner. As shown in Table 9, 2-stage k-means leads to marginally worse general QA performance (-0.37 VQAv2), at the same time improves OCR performance (+0.78 TextVQA).
Method VQAv2 TextVQA MME RefCOCO
val testA testB
balanced k-means 72.19 50.91 1475.78 75.47 80.20 71.36
2-stage balanced k-means 71.82 51.69 1421.01 74.76 80.99 69.99
Table 9: The impact of clustering algorithm
6 Conclusion
In this work, we present a theoretical analysis of decentralization of autoregressive generation. We show that AR sampling is a special case of Discrete Flow Matching in discrete time domain. We also define Discrete Decentralized Flow Matching objective, by expressing probability generating velocity as a linear combination of expert flows. We also conduct experiments, showing the parity between decentralized and centralized training settings for MLLMs across diverse set of benchmarks.
The possible future directions include finding new more optimal dataset partitioning algorithms, scaling the model sizes and datasets, and finding a unified routing strategy for both image and text samples.

### v2

@@
 1 Introduction
-The increase in performance of Multimodal Large Language Models (MLLMs) requires a massive increase in parameter counts (kaplan2020scalinglawsneurallanguage), which, in turn, requires enhanced infrastructure: training requires centralized clusters of thousands of GPUs with extensive gradient synchronization.
+The increase in performance of Multimodal Large Language Models (MLLMs) requires a massive increase in parameter counts (Kaplan et al., 2020), which, in turn, requires enhanced infrastructure: training requires centralized clusters of thousands of GPUs with extensive gradient synchronization.
 These developments present two critical challenges.
 First, training LLMs is inaccessible to academic institutions and small startups due to limited infrastructure.
-Second, the system is susceptible to individual node failures, which can disrupt the entire process (grattafiori2024llama3herdmodels).
+Second, the system is susceptible to individual node failures, which can disrupt the entire process (Grattafiori et al., 2024).
 Existing mitigation strategies include aggressive checkpointing and related techniques.
@@
 This significantly reduces communication between nodes and enables academic institutions to collaborate.
-Recent studies (li2022branchtrainmergeembarrassinglyparalleltraining; gururangan2023scalingexpertlanguagemodels; no-need-to-talk) empirically demonstrate the efficiency of this approach for unimodal language modeling training in specific settings; however, existing work lacks a strict theoretical guarantee that decentralized training is equivalent to the centralized paradigm, making decentralized training a black box.
+Recent studies (Li et al., 2022; Gururangan et al., 2023; Filippova et al., 2025) empirically demonstrate the efficiency of this approach for unimodal language modeling training in specific settings; however, existing work lacks a strict theoretical guarantee that decentralized training is equivalent to the centralized paradigm, making decentralized training a black box.
 Moreover, decentralized training of multimodal large language models has not yet been empirically explored.
-mcallister2025decentralizeddiffusionmodels defined the Decentralized Flow Matching objective for Continuous Flow Matching (lipman2023flow) and demonstrated the equivalence between centralized and decentralized training of flow-matching models for image generation.
-At the same time, gat2024discrete introduced the Discrete Flow Matching theoretical framework, extending Continuous Flow Matching (lipman2023flow) to a discrete data domain.
+McAllister et al.
+(2025) defined the Decentralized Flow Matching objective for Continuous Flow Matching (Lipman et al., 2023) and demonstrated the equivalence between centralized and decentralized training of flow-matching models for image generation.
+At the same time, Gat et al.
+(2024) introduced the Discrete Flow Matching theoretical framework, extending Continuous Flow Matching (Lipman et al., 2023) to a discrete data domain.
 Inspired by these advancements, we derive the theoretical analysis of decentralized autoregressive generation.
 Our contributions are the following:
-  • We extend the Discrete Flow Matching paradigm (gat2024discrete) to the discrete time domain, making the necessary modifications and defining the corresponding conditions.
-  • Similar to mcallister2025decentralizeddiffusionmodels, we define the Decentralized Discrete Flow Matching objective and show that the probability-generating velocity is a linear combination of expert flows.
+  • We extend the Discrete Flow Matching paradigm (Gat et al., 2024) to the discrete time domain, making the necessary modifications and defining the corresponding conditions.
+  • Similar to McAllister et al. (2025), we define the Decentralized Discrete Flow Matching objective and show that the probability-generating velocity is a linear combination of expert flows.
   • We show that autoregressive generation is a special degenerate case of the Discrete Flow Matching theoretical paradigm.
 We present the corresponding probability path and probability generating velocity, implying that the Decentralized Discrete Flow Matching objective can be applied to autoregressive generation.
-  • We conduct experiments demonstrating the equivalence between decentralized and centralized training for multimodal language models in two distinct settings: LLaVA (liu2024improvedbaselinesvisualinstruction) and InternVL (chen2025expandingperformanceboundariesopensource).
+  • We conduct experiments demonstrating the equivalence between decentralized and centralized training for multimodal language models in two distinct settings: LLaVA (Liu et al., 2024) and InternVL (Chen et al., 2025).
 2 Background
@@
 2.1 Continuous Time Discrete Flow Matching
-Discrete Flow Matching (gat2024discrete) is a theoretical framework that describes discrete generative modeling processes in terms similar to Continuous Flow Matching (lipman2023flow).
+Discrete Flow Matching (Gat et al., 2024) is a theoretical framework that describes discrete generative modeling processes in terms similar to Continuous Flow Matching (Lipman et al., 2023).
 Namely, having vocabulary or tokens $[d]$, source distribution $X_{0}$ with PMF $p(x)$ of token sequences $x=(x^{1},\ldots,x^{N})\in[d]^{N}$ and target distribution $X_{1}$ with PMF $q(x)$, we define the probability path $p_{t}(x),t\in[0,1]$ s.t.
@@
 where $\delta_{z}(x^{\bar{i}})=\prod_{j\neq i}\delta_{z}(x^{j})$.
-As shown in (gat2024discrete), if the continuity equation holds, then the probability generating velocity generates a probability path as defined in (7-8).
+As shown in (Gat et al., 2024), if the continuity equation holds, then the probability generating velocity generates a probability path as defined in (7-8).
 3 Theoretical Analysis
@@
 3.2 Autoregressive sampling
-To define autoregressive sampling, we define the source and target distribution as a special case of C-coupling from (gat2024discrete):
+To define autoregressive sampling, we define the source and target distribution as a special case of C-coupling from (Gat et al., 2024):
 $(X_{0},X_{1})\sim({\mathbb{I}}\odot X_{1}+(1-{\mathbb{I}})\odot(\mathbf{m},\mathbf{m},\ldots,\mathbf{m}),X_{1})$ (18)
@@
 The above formalization automatically makes Theorem 1 true for autoregressive sampling.
-Now, similar to (mcallister2025decentralizeddiffusionmodels), if we define a target distribution $X_{1}$ as a union of disjoint clusters $S_{k},\quad k\in\{1,2,\dots,K\}$, we can rewrite Equation 24 as follows:
+Now, similar to (McAllister et al., 2025), if we define a target distribution $X_{1}$ as a union of disjoint clusters $S_{k},\quad k\in\{1,2,\dots,K\}$, we can rewrite Equation 24 as follows:
 $u_{t}^{i}(x^{i},z)=\sum_{(x_{0},x_{1})\in[d]^{N\times N}}u_{t}^{i}(x^{i},z|x_{0},x_{1})\frac{p_{t}(z|x_{0},x_{1})\pi(x_{0},x_{1})}{p_{t}(z)}$ (24)
@@
 4 Method
-Since the effectiveness of decentralized discrete flow matching in case of pure LM has already been shown in previous work (no-need-to-talk; gururangan2023scalingexpertlanguagemodels), we focus on Vision Language Models as another instance of autoregressive sampling.
+Since the effectiveness of decentralized discrete flow matching in case of pure LM has already been shown in previous work (Filippova et al., 2025; Gururangan et al., 2023), we focus on Vision Language Models as another instance of autoregressive sampling.
 4.1 Expert training
@@
 5.1 LLaVA
-In the first part of the experiments, we utilize the LLaVA-1.5 (liu2024improvedbaselinesvisualinstruction) fine-tuning dataset, consisting of 625k image-text samples and 41k text-only samples from diverse sources.
+In the first part of the experiments, we utilize the LLaVA-1.5 (Liu et al., 2024) fine-tuning dataset, consisting of 625k image-text samples and 41k text-only samples from diverse sources.
 During the data partitioning stage, text-only samples are randomly and equally distributed between the clusters.
@@
 5.2 InternVL
-In the second phase of our experiments, we utilize a subset of the InternVL 2.5 (chen2025expandingperformanceboundariesopensource) Stage 2 fine-tuning data mixture (see Table 1).
+In the second phase of our experiments, we utilize a subset of the InternVL 2.5 (Chen et al., 2025) Stage 2 fine-tuning data mixture (see Table 1).
 We curated datasets representing diverse task types to ensure comprehensive coverage.
@@
 Task type Datasets
-Captioning TextCaps (sidorov2020textcapsdatasetimagecaptioning) , ShareGPT4o (cui2025comprehensive)
-General QA VQAv2(balanced_vqa_v2), GQA(hudson2019gqanewdatasetrealworld)
-Chart ChartQA(masry2022chartqabenchmarkquestionanswering)
-OCR InfoVQA(mathew2021infographicvqa) , TextVQA(singh2019vqamodelsread)
-Knowledge KVQA(shahMYP19), A-OKVQA(schwenk2022aokvqabenchmarkvisualquestion)
-Document DocVQA(mathew2021docvqadatasetvqadocument)
-Grounding RefCOCO/+/g (mao2016generationcomprehensionunambiguousobject; yu2016modelingcontextreferringexpressions)
-Science AI2DKembhavi2016ADI, ScienceQAlu2022learn
-Conversation ALLaVA(chen2024allavaharnessinggpt4vsynthesizeddata)
+Captioning TextCaps (Sidorov et al., 2020) , ShareGPT4o (Cui et al., 2024)
+General QA VQAv2(Goyal et al., 2017), GQA(Hudson and Manning, 2019)
+Chart ChartQA(Masry et al., 2022)
+OCR InfoVQA(Mathew et al., 2021a) , TextVQA(Singh et al., 2019)
+Knowledge KVQA(Sanket Shah and Talukdar, 2019), A-OKVQA(Schwenk et al., 2022)
+Document DocVQA(Mathew et al., 2021b)
+Grounding RefCOCO/+/g (Mao et al., 2016; Yu et al., 2016)
+Science AI2DKembhavi et al. (2016), ScienceQALu et al. (2022)
+Conversation ALLaVA(Chen et al., 2024)
 Table 1: InternVL-type models fine-tuning data mixture
 5.3 Evaluation
-We evaluate our LLaVA models across total of 8 benchmarks, used in the liu2024improvedbaselinesvisualinstruction.
-For InternVL, we evaluate across 14 benchmarks, aiming to assess the models in the general question answering, OCR and other tasks, used in the chen2025expandingperformanceboundariesopensource.
+We evaluate our LLaVA models across total of 8 benchmarks, used in the Liu et al.
+(2024).
+For InternVL, we evaluate across 14 benchmarks, aiming to assess the models in the general question answering, OCR and other tasks, used in the Chen et al.
+(2025).
 5.4 Results
@@
 In the main InternVL experiments, we use spherical balanced k-means clustering algorithm for dataset partitioning and routing.
-To test the impact of the clustering algorithm, we also train $K=2$ experts using 2-stage balanced spherical k-means algorithm inspired by mcallister2025decentralizeddiffusionmodels.
+To test the impact of the clustering algorithm, we also train $K=2$ experts using 2-stage balanced spherical k-means algorithm inspired by McAllister et al.
+(2025).
 Namely, first stage is fine clustering into $k=1024$ clusters in unbalanced manner.